OpenStudy (anonymous):

How would you convert the repeating, non terminating decimal into a fraction? Explain the process as you solve the problem 0.1515.

OpenStudy (anonymous):

\[0.1515...=\frac{a}{b}\\ 15.1515...=\frac{100a}{b}\] Subtract the first equation from the second: \[15=\frac{99a}{b}~~\iff~~\frac{a}{b}=\frac{15}{99}\]